8. A train travels 360 km at a uniform speed. If the speed had been 5 km/h more, it would have taken 1 hour less for the same journey. Find the speed of the train.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let us say, the speed of the train be x km/hr. Time taken to cover 360 km = 360/x hr. As per the question given, ⇒ (x + 5)(360-1/x) = 360 ⇒ 360 – x + 1800-5/x = 360 ⇒ x2 + 5x + 10x – 1800 = 0 ⇒ x(x + 45) -40(x + 45) = 0 ⇒ (x + 45)(x – 40) = 0 ⇒ x = 40, -45 As we know, the value of speed cannot be negative. Therefore, the speed of train is 40 km/h.